# FlatCAD: Fast Curvature Regularization of Neural SDFs for CAD Models

### v1

@@ Abstract.
Neural signed-distance fields (SDFs) have become a versatile backbone for geometric learning, yet enforcing developable, CAD-style behavior still hinges on Gaussian-curvature penalties that require full Hessian evaluation and second-order automatic differentiation, which are both costly in memory and runtime. We present a curvature proxy that regularities only the mixed second-order term (Weingarten term), allowing the two principal curvatures to adapt freely to data while suppressing unwanted warp. Two complementary instantiations realize this idea: (i) a finite-difference proxy that replaces every Hessian entry with four forward SDF evaluations and a single first-order gradient, and (ii) an autodiff proxy that arrives at the same mixed derivative via one Hessian–vector product, sidestepping explicit full Hessian assembly remaining faster in practice. Both variants converge to the exact mixed second derivative, thus preserving the intended geometric bias without incurring full second-order graphs. On the ABC benchmarks the proxies match or exceed the reconstruction fidelity of Hessian-based baselines while cutting GPU memory and wall-clock time by 2×. Because the method is drop-in and framework-agnostic, it opens a practical path toward scalable, curvature-aware SDF learning for engineering-grade shape reconstruction.
@@
copyright: rightsretainedccs: Computing methodologies Computer graphicsccs: Computing methodologies Neural networksccs: Computing methodologies Computational geometry and object modelingccs: Computing methodologies Shape modeling
Figure 1. Comparison of our Curvature Proxy (top row) and the Full-Hessian-based method (bottom row) on a CAD part reconstruction: both achieve similar developable‐quality surfaces, but our approach converges in roughly half the time.
@@ 1. Introduction
However, for accurate geometric representations, a remaining challenge was geometric faithfulness: vanilla SDF fits can satisfy point-wise losses while remaining wildly inconsistent in surface detail or curvature. Lipman et al. addressed this with the Implicit Geometric Regularizer (IGR) (Gropp et al., 2020), adding an Eikonal term $\lVert\nabla f\rVert\approx 1$ that enforces unit-length gradients, followed by second-order methods like DiGS (Ben-Shabat et al., 2021), which stabilizes training using the divergence, and Neural Singular Hessian (Wang et al., 2023), which pushes the network’s Hessian toward low rank, which discourages spurious curvature and enforces piece-wise smoothness that meshes better with classical differential geometry.
@@
That thread naturally intersects with computer-aided design (CAD). Mechanical parts are dominated by plates, cylinders, cones, and developable blends whose Gaussian curvature is zero almost everywhere, save for sharp feature curves. Encoding this developability in an SDF prior promises more faithful reconstructions from sparse or noisy scans and cleaner patch decomposition for downstream parametric recovery. NeurCADRecon (Dong et al., 2024) recently embodied that idea: it augments a Siren with a “developability” loss that directly integrates the magnitude of Gaussian curvature $|K|$ over a thin shell around the surface, thereby encouraging each learned patch to flatten out unless data dictates otherwise.
@@
Unfortunately, such elegance comes at a price. The Gaussian curvature of an implicit surface depends on all nine entries of the Hessian of the network. In automatic differentiation frameworks this means six Jacobian-vector products—and thus seven backward passes—per sample, plus the retention of second-order computational graphs. Memory footprints balloon, forcing tiny batches; iteration times rise, limiting practical resolutions.
Our Contributions
@@
To address these limitations we propose a novel curvature regularization loss which we call curvature proxy. In partcular, the contributions of the paper are the following:
@@
  • Curvature proxy. We regularize only the mixed Weingarten term, letting each principal curvature follow the data while flattering hyperbolic and parabolic patches and rounds elliptic ones uniformly.
@@
  • Two implementations. We propose two variants for computing the proxy. (i) Finite-differences, which uses four additional forward SDF queries plus one gradient evaluation and introduces no second-order graphs; (ii) Auto-diff proxy, which employs a single Hessian–vector product obtained with two reverse sweeps, thereby avoiding full-Hessian assembly and running faster in practice. In both cases the estimate converges to the exact mixed derivative as the tangent step shrinks.
@@
  • Efficiency without loss of accuracy. On the ABC CAD benchmark our proxies match—or slightly improve—Hessian‐based baselines while roughly halving GPU memory use and training time.
@@
Because the method is drop‐in and framework‐agnostic, it makes curvature‐aware neural SDFs practical for large‐scale, engineering‐grade shape reconstruction.
@@ 2. Related Work
Early implicit methods compute the signed distance to the tangent plane of the closest surface point (Hoppe et al., 1992). Subsequently, radial basis function (RBF) methods were introduced to fit the zero level-set of the signed distance function (SDF), enabling smooth surface reconstruction (Carr et al., 2001; Huang et al., 2019; Li et al., 2016). Another family of methods formulates reconstruction as solving partial differential equations, such as Poisson Surface Reconstruction (PSR) (Kazhdan et al., 2020; Kazhdan and Hoppe, 2013; Kazhdan et al., 2006; Hou et al., 2022), which computes an occupancy field by solving a Poisson equation. Parametric Gauss Reconstruction (PGR) (Lin et al., 2022) improves normal consistency by leveraging the Gauss formula from potential theory. These methods build surfaces analytically by imposing strong geometric constraints.
@@
To further enhance geometric fidelity, higher-order constraints such as curvature have also been explored. The Hessian matrix of a neural field encodes its second-order derivatives, making it a natural vehicle for expressing differential properties such as surface curvature. DiGS (Ben-Shabat et al., 2021) penalizes the divergence of the gradient field as a soft constraint, encouraging smooth transitions across the surface. Neural-Singular-Hessian (Wang et al., 2023) directly regularizes the Hessian by minimizing the smallest singular value of the Hessian to make it rank-deficient near the surface, promoting developability and reducing spurious oscillations. These second-order constraints have proven effective in restoring fine-grained geometric structures. Concurrently, other curvature-guided regularizers have been explored for surface INRs (Sang et al., 2025; Novello et al., 2023), and rank-minimization schemes have been proposed to approximate developability (Selvaraju, 2024).
@@
Figure 2. Principal curvatures $\kappa_{1}$ and $\kappa_{2}$ are the eigen-values of the Weingarten map $S$. Their product give the Gaussian curvature $K=\kappa_{1}\kappa_{2}=\det S$. Because $S$ is self-adjoint with respect to the first fundamental form, its eigen-vectors (the principal directions) are orthogonal in the surface metric whenever the eigen-values are distinct. At an umbilic point (e.g. on a sphere) the entire two-dimensional tangent plane is the eigen-space, so no unique pair of principal directions exists. For a plane the second fundamental form vanishes, i.e., $S=0$.
@@ 3. Method
Our goal is to learn a neural signed-distance field $f$ from an unoriented point cloud such that its zero-set reproduces the target CAD surface. We keep the standard supervision recipe used by implicit reconstruction methods and add a minimal curvature prior that removes local warp, flattening developable and saddle parts and rounding elliptic ones–without ever building a full Hessian.
@@
Second, and most importantly, we introduce a novel curvature proxy that suppresses the difference between the principal curvatures of the learned surface. This regularization is particularly effective for CAD surfaces, as it forces parabolic and hyperbolic surface parts to be planar and elliptic areas to be umbilical (cf. Figure 2). In combination with the surface reconstruction and Eikonal losses, promotes flatness where the target is hyperbolic and parabolic, and spherical uniformity where it is doubly curved.
Unlike previous methods that construct the full $3\times 3$ Hessian(Wang et al., 2023; Dong et al., 2024), our formulation needs only a single Hessian–vector product per sample, eliminating explicit Hessian assembly and speeding training. Additionally, we also propose a first order finite-difference approximation, which avoids Hessian computation entirely, enabling significantly faster training while retaining the geometric benefits of curvature regularization.
3.1. Background and Motivation
Signed-Distance Fields and Surface Geometry
We represent a surface $\mathcal{S}\subset\mathbb{R}^{3}$ implicitly as the zero-level set of a signed-distance field (SDF)
$f:\mathbb{R}^{3}\!\longrightarrow\!\mathbb{R},\qquad\mathcal{S}=\bigl{\{}\mathbf{x}\mid f(\mathbf{x})=0\bigr{\}}.$
Because a true SDF satisfies $\lVert\nabla f\rVert=1$ in a neighbourhood of $\mathcal{S}$, its normal direction is locally linear:
@@ 3.1. Background and Motivation
That operator (Weingarten map) takes any tangent vector and returns how fast the surface normal rotates in that direction, i.e. the local bending. Its eigenvalues are the principal curvatures $\kappa_{1},\kappa_{2}$; their product gives the Gaussian curvature $K=\kappa_{1}\kappa_{2}=\det S$, whose sign distinguishes elliptic, parabolic and hyperbolic regimes as illustrated in Fig. 2.
@@
In the principal frame $S=\operatorname{diag}(\kappa_{1},\kappa_{2})$, rotating the basis by an angle $\theta$ in the tangent plane gives
$S(\theta)=\begin{pmatrix}\kappa_{1}\cos^{2}\!\theta+\kappa_{2}\sin^{2}\!\theta&(\kappa_{2}-\kappa_{1})\sin\theta\cos\theta\\[4.0pt] (\kappa_{2}-\kappa_{1})\sin\theta\cos\theta&\kappa_{1}\sin^{2}\!\theta+\kappa_{2}\cos^{2}\!\theta\end{pmatrix}.$
@@
Expected Effect
During training we penalize $\lvert S_{12}\rvert$, in particular, for each point we sample a fresh random frame angle $\theta$ each iteration. Taking the expectation of $\lvert S_{12}\rvert^{2}$ over $\theta\in[0,2\pi)$ yields
(2) $\mathbb{E}_{\theta}\bigl{[}S_{12}^{2}\bigr{]}=\frac{(\kappa_{2}-\kappa_{1})^{2}}{4}\,\mathbb{E}_{\theta}\!\bigl{[}\sin^{2}2\theta\bigr{]}=\frac{(\kappa_{2}-\kappa_{1})^{2}}{8}.$
Equation (2) vanishes iff $\kappa_{1}=\kappa_{2}$. Hence gradient descent on $\lvert S_{12}\rvert$ systematically suppresses the curvature gap without directly pushing either curvature to zero.
Because the frame angle is resampled independently as $\theta\sim\mathcal{U}[0,2\pi)$, the off-diagonal entry $S_{12}$ is a zero-mean random variable. The mini-batch average of $S_{12}^{2}$ (or $|S_{12}|$) is thus an unbiased Monte-Carlo estimate of its expectation over $\theta$. If the measurement noise on $S_{12}(\theta)$ is assumed Gaussian with variance $\sigma^{2}$, the negative log-likelihood reduces—up to a constant—to
@@
Minimizing the squared proxy is therefore the maximum-likelihood estimator for the zero-warp hypothesis. Adopting a Laplace noise model instead yields the absolute-value proxy, with the same Monte-Carlo interpretation.
Hessian Deficiency
Evaluating $S_{12}$ directly on the zero level-set is numerically unstable: the Hessian is rank-deficient because $\partial_{nn}f=0$. To address it we follow the sampling strategy of Dong et al. (Dong et al., 2024): At each iteration, we sample a static shell $\Omega$ by drawing one point around every input point $p$ from a 3D Gaussian whose standard deviation equals the distance to its 50-th nearest neighbor, yielding a fixed batch of 15 k near-surface points. We evaluate $S_{12}$ on $\Omega$, which provides stable curvature signals while leaving the standard data and Eikonal terms unchanged.
The off-diagonal proxy is hence a minimal, orientation-free condition that suppresses the curvature difference $(\kappa_{2}-\kappa_{1})^{2}$, thereby isolating the spurious “warp” defect that is not part of the many intended CAD shapes, and flattening hyperbolic and parabolic areas. It is evaluated with a single Hessian-vector product $H_{f}\,\mathbf{v}$, avoiding full Hessian construction. Because the loss depends only linearly on this mixed second derivative, rather than on quadratic combinations of Hessian entries, it yields well-conditioned, low-variance gradients that integrate seamlessly with standard SDF and Eikonal losses, mitigating the exploding sensitivities observed with many quadratic-or-higher curvature regularizes.

### v2

@@ Abstract.
Neural signed-distance fields (SDFs) are a versatile backbone for neural geometry representation, but enforcing CAD-style developability usually requires Gaussian-curvature penalties with full Hessian evaluation and second-order differentiation, which are costly in memory and time. We introduce an off-diagonal Weingarten loss that regularizes only the mixed shape operator term that represents the gap between principal curvatures and flattens the surface. We present two variants: a finite-difference version using six SDF evaluations plus one gradient, and an auto-diff version using a single Hessian-vector product. Both converge to the exact mixed term and preserve the intended geometric properties without assembling the full Hessian. On the ABC benchmarks the losses match or exceed Hessian-based baselines while cutting GPU memory and training time by roughly a factor of two. The method is drop-in and framework-agnostic, enabling scalable curvature-aware SDF learning for engineering-grade shape reconstruction. Our code is available at https://flatcad.github.io/.
@@
copyright: rightsretainedccs: Computing methodologies Computer graphicsccs: Computing methodologies Shape modelingccs: Computing methodologies Regularizationccs: Computing methodologies Machine learning algorithms
Figure 1. Comparison of our Weingarten regularizer (top row) and the full-Hessian-based method of previous work (Dong et al., 2024) (bottom row) on a CAD part reconstruction: both achieve similar-quality surfaces, but our method converges in roughly half the time and uses about half the memory.
@@ 1. Introduction
However, for accurate geometric representations, a remaining challenge was geometric faithfulness: vanilla SDF fits can satisfy point-wise losses while remaining wildly inconsistent in surface detail or curvature. Lipman et al. addressed this with the Implicit Geometric Regularizer (IGR) (Gropp et al., 2020), adding an Eikonal term $\lVert\nabla f\rVert\approx 1$ that enforces unit-length gradients, followed by second-order methods like DiGS (Ben-Shabat et al., 2021), which stabilizes training using the divergence, and Neural-Singular-Hessian (Wang et al., 2023), which pushes the network’s Hessian toward low rank, which discourages spurious curvature and enforces piece-wise smoothness that matches better with classical differential geometry.
@@
That thread naturally intersects with computer-aided design (CAD). Mechanical parts are dominated by plates, cylinders, cones, and developable blends whose Gaussian curvature is zero almost everywhere, except for sharp feature curves. Encoding these properties in an SDF prior promises more faithful reconstructions from sparse or noisy scans and cleaner patch decomposition for downstream parametric recovery. NeurCADRecon (Dong et al., 2024) recently embodied that idea: it augments a SIREN network with a “developability” loss that minimizes the magnitude of Gaussian curvature $|K|$ over a thin shell around the surface, thereby encouraging each learned patch to flatten out unless data dictates otherwise.
@@
Unfortunately, such elegance comes at a price. The Gaussian curvature of an implicit surface depends on all nine entries of the Hessian of the network. In automatic differentiation frameworks this means six Jacobian-vector products and thus seven backward passes per sample, plus the retention of second-order computational graphs. Memory footprints balloon, forcing small batches; iteration times rise, limiting practical resolutions.
@@
To address these limitations we propose a novel curvature regularization loss. In particular, our contributions are the following:
@@
  • Curvature gap regularization. We regularize only the off-diagonal Weingarten shape operator term, letting each principal curvature follow the data while flattening hyperbolic and parabolic patches and rounding elliptic ones uniformly. In other words, we minimize the “curvature-gap” which is the difference between the principal curvatures and we denote the resulting regularizer the off-diagonal Weingarten loss (ODW-loss). It is described in detail in Section 3.
@@
  • Two implementations. We propose two variants for computing the regularizer: (i) Finite‐differences, which uses forward and backward offsets (six additional SDF queries plus one gradient evaluation) and introduces no second-order graphs; (ii) Auto-differentiation, which employs a single Hessian-vector product obtained with two reverse sweeps, thereby avoiding full-Hessian assembly and running faster in practice. Both are described in detail in Section 4.
@@
  • Efficiency without loss of accuracy. On the ABC CAD benchmark our proxies match or slightly outperform Hessian-based baselines while roughly halving GPU memory use and training time. The details are presented in Section 5.
@@
Because the method is drop‐in and framework‐agnostic, it makes curvature‐aware neural SDFs practical for large‐scale, engineering‐grade shape reconstruction. Our implementation is available at https://flatcad.github.io/.
@@ 2. Related Work
Early implicit methods compute the signed distance to the tangent plane of the closest surface point (Hoppe et al., 1992). Subsequently, radial basis function (RBF) methods were introduced to fit the zero level-set of the signed distance function, enabling smooth surface reconstruction (Carr et al., 2001; Huang et al., 2019; Li et al., 2016). Another family of methods formulates reconstruction as solving partial differential equations, such as Poisson Surface Reconstruction (PSR) (Kazhdan et al., 2020; Kazhdan and Hoppe, 2013; Kazhdan et al., 2006; Hou et al., 2022), which computes an occupancy field by solving a Poisson equation. Parametric Gauss Reconstruction (PGR) (Lin et al., 2022) improves normal consistency by leveraging the Gauss formula from potential theory. These methods build surfaces analytically by imposing strong geometric constraints.
@@
To further enhance geometric fidelity, higher-order constraints have also been explored. The Hessian matrix of a neural field encodes its second-order derivatives, making it a natural vehicle for expressing differential properties such as surface curvature. DiGS (Ben-Shabat et al., 2021) penalizes the divergence of the gradient field as a soft constraint, encouraging smooth transitions across the surface. Neural-Singular-Hessian (Wang et al., 2023) directly regularizes the Hessian by minimizing the smallest singular value to make it rank-deficient near the surface, promoting developability and reducing spurious oscillations. These second-order constraints have proven effective in restoring fine-grained geometric structures. Concurrently, other curvature-guided regularizers have been explored for surface INRs (Sang et al., 2025; Novello et al., 2023), and rank-minimization schemes have been proposed to approximate developability (Selvaraju, 2024).
@@
Figure 2. Principal curvatures $\kappa_{1}$ and $\kappa_{2}$ are the eigenvalues of the Weingarten map $S$. Their product gives the Gaussian curvature $K=\kappa_{1}\kappa_{2}=\det S$. Because $S$ is self-adjoint with respect to the first fundamental form, its eigenvectors (the principal directions) are orthogonal in the surface metric whenever the eigenvalues are distinct. At an umbilic point (e.g. on a sphere) the entire two-dimensional tangent plane is the eigenspace, so no unique pair of principal directions exists. For a plane the second fundamental form vanishes, i.e., $S=0$.
@@ 3. Method
Our goal is to learn a neural signed-distance field $f$ from an unoriented point cloud such that its zero-set reproduces the target CAD surface. We keep the standard supervision recipe used by implicit reconstruction methods and add a minimal curvature regularizer that removes local warp, flattening developable and saddle parts and rounding elliptic ones–without ever building a full Hessian.
@@
Second, and most importantly, we introduce a novel curvature regularizer that suppresses the difference between the principal curvatures of the learned surface. This regularization is particularly effective for CAD surfaces, as it forces parabolic and hyperbolic surface parts to be planar and elliptic areas to be umbilical (cf. Figure 2). In combination with the surface reconstruction (Dirichlet) and unit gradient field (Eikonal) losses, it promotes flatness where the target is hyperbolic and parabolic, and spherical uniformity where it is doubly curved.
Unlike previous methods that construct the full $3\times 3$ Hessian (Wang et al., 2023; Dong et al., 2024), our formulation needs only a single Hessian–vector product per sample, eliminating explicit Hessian assembly and speeding up training. Additionally, we also propose a first-order finite-difference approximation, which avoids Hessian computation entirely, enabling significantly faster training while retaining the geometric benefits of curvature regularization.
3.1. Signed-Distance Fields and Surface Geometry
We represent a surface $\mathcal{M}\subset\mathbb{R}^{3}$ implicitly as the zero-level set of a signed-distance field (SDF)
$f:\mathbb{R}^{3}\!\longrightarrow\!\mathbb{R},\qquad\mathcal{M}=\bigl{\{}\mathbf{x}\mid f(\mathbf{x})=0\bigr{\}}.$
Because a true SDF satisfies $\lVert\nabla f\rVert=1$ in a neighborhood of $\mathcal{M}$, its normal direction is locally linear:
@@ 3.1. Signed-Distance Fields and Surface Geometry
That operator (Weingarten map) takes any tangent vector and returns how fast the surface normal rotates in that direction, i.e., the local bending. Its eigenvalues are the principal curvatures $\kappa_{1},\kappa_{2}$; their product gives the Gaussian curvature $K=\kappa_{1}\kappa_{2}=\det S$. Together, these values distinguish spherical, planar, parabolic, elliptic, and hyperbolic regimes as illustrated in Fig. 2.
@@
In the principal frame the shape operator is
$S_{0}=\begin{pmatrix}\kappa_{1}&0\\[2.0pt] 0&\kappa_{2}\end{pmatrix}$.
We can rotate a tangent basis by an angle $\theta$ through
The operator in the rotated frame is the orthogonal transform
@@
3.2. Maximum Likelihood Estimation
During training we penalize $S_{12}^{2}$ (or $|S_{12}|$, see below), in particular, for each point we sample a fresh random frame angle $\theta$ each iteration (which in practice is equivalent to sampling a fresh random vector $\mathbf{u}$ orthogonal to $\mathbf{n}$, and completing the frame with $\mathbf{v}$ using the cross product).
Taking the expectation of $S_{12}^{2}$ over $\theta\in[0,2\pi)$ yields
Because
we obtain
(2) $\mathbb{E}_{\theta}[S_{12}^{2}]=\frac{1}{8}(\kappa_{2}-\kappa_{1})^{2}.$
Equation (2) vanishes iff $\kappa_{1}=\kappa_{2}$. Hence gradient descent on $S_{12}^{2}$ systematically suppresses the curvature gap without directly pushing either curvature to zero.
Because the frame angle is resampled independently as $\theta\sim\mathcal{U}[0,2\pi)$, the off-diagonal Weingarten entry $S_{12}$ is a zero-mean random variable. The mini-batch average of $S_{12}^{2}$ (or $|S_{12}|$) is thus an unbiased Monte-Carlo estimate of its expectation over $\theta$. If the measurement noise on $S_{12}(\theta)$ is assumed Gaussian with variance $\sigma^{2}$, the negative log-likelihood reduces—up to a constant—to
@@ 3.2. Maximum Likelihood Estimation
Minimizing the squared off-diagonal entry is therefore the maximum-likelihood estimator for the zero-warp hypothesis. Consequently, the $L^{2}$ formulation possesses a clear maximum-likelihood interpretation under the common Gaussian-noise assumption.
Expectation of the absolute value
Adopting a Laplace noise model instead yields the absolute-value, with the same Monte-Carlo interpretation. With $\theta\sim\mathrm{Uniform}[0,2\pi)$,
$\,=\,\lvert\kappa_{2}-\kappa_{1}\rvert\;\frac{1}{2\pi}\int_{0}^{2\pi}\!\lvert\sin\theta\cos\theta\rvert\,d\theta.$
Evaluating the integral:
$\int_{0}^{2\pi}\!\lvert\sin\theta\cos\theta\rvert\,d\theta=\frac{1}{2}\int_{0}^{2\pi}\!\lvert\sin 2\theta\rvert\,d\theta=\frac{1}{2}\!\times\!4=2$
yields
(3) $\mathbb{E}_{\theta}\bigl{[}\lvert S_{12}\rvert\bigr{]}=\frac{1}{\pi}\lvert\kappa_{2}-\kappa_{1}\rvert$.
While the squared Weingarten loss has elegant Gaussian-noise and variance interpretations, in a SIREN setting the practical stability, edge preservation, and faster optimization of the absolute-value off-diagonal Weingarten loss generally outweigh the theoretical neatness of $L^{2}$.
SIREN network’s high-frequency sine layers can generate very large local gradients; employing the absolute ($L^{1}$) loss caps each sample’s gradient at $\pm 1$, preventing exploding updates while preserving sharp detail, and thus trains more stably than the squared ($L^{2}$) version.
3.3. Hessian Deficiency
Evaluating $S_{12}$ directly on the zero level-set is numerically unstable: the Hessian is rank-deficient because $\partial_{nn}f=0$. In fact, since the Eikonal condition enforces $\|\nabla f\|=1$, differentiating this constraint along the normal direction yields
$\partial_{n}\!\left(\tfrac{1}{2}\|\nabla f\|^{2}\right)\;=\;(H_{f}\mathbf{n})\cdot\nabla f\;=\;\mathbf{n}^{\top}H_{f}\mathbf{n}\;=\;0,$
which means that one eigenvalue of $H_{f}$ must vanish, making the Hessian singular on the surface. This property of signed distance fields is also emphasized by Wang et al. (Wang et al., 2023), who use it as the basis for their Neural-Singular-Hessian regularizer.
To address the numerical instability, we follow the sampling strategy of Dong et al. (Dong et al., 2024). At each iteration, we sample a static shell $\Omega$ by drawing one point around every input point $p$ from a 3D Gaussian whose standard deviation equals the distance to its 50-th nearest neighbor, yielding a fixed batch of 15k near-surface points. We then evaluate $S_{12}$ on $\Omega$, which provides stable signals while leaving the standard data and Eikonal terms unchanged.
The off-diagonal Weingarten loss is hence a minimal, orientation-free condition that suppresses the curvature difference $(\kappa_{2}-\kappa_{1})^{2}$, thereby isolating the spurious “warp” defect and flattening hyperbolic and parabolic areas.
Because the loss depends only linearly on this mixed second derivative, rather than on quadratic combinations of Hessian entries, it yields well-conditioned, low-variance gradients that integrate seamlessly with standard SDF and Eikonal losses, mitigating the exploding sensitivities observed with many quadratic-or-higher curvature regularizers.